OpenStudy (anonymous):

One body of water covers approximately 70 million square miles. Another covers approximately 36 million square miles. What percent of the area covered by the first sea is the area covered by the second? Round to the nearest hundredth.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

as far as i know, the difference x 100/intial value (70-36)x100/70=24x100/70 = 34.285%

OpenStudy (anonymous):

Thankyou

OpenStudy (anonymous):

i am not 100% sure though xD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

meh i fink, its 36x100/70= 51.428%

OpenStudy (anonymous):

ok